### v5

@@ V.2 Exploration of the topological invariance
In Appendix B, we have explained how the SRG is theoretically related to the persistent homology theory and other concepts, where we suggest the potential applicability of the SRG in analyzing persistent homology. Below, we validate this applicability by using the SRG to explore the topological invariance.
In brief, the persistent homology theory can be used to study the lifetimes of different topological properties in a filtration process, during which we progressively add or remove simplices from the system [62, 74, 75, 76] (i.e., this is similar to the network evolution process in complex network studies [82]). Although the SRG is not exactly same as the classic filtration approach because it modifies the system in a more complicated manner, it does offer an opportunity to study the evolution of topological properties during renormalization (i.e., we can measure different topological properties on $\mathbf{G}^{\left(p\right)}_{k}$ for each $k\geq 1$). Similar to the analysis of scale-invariance, a topological property can be treated as invariant if it is persistent across all the iterations of the SRG (i.e., referred to as the topological invariance). In our analysis, we mainly focus on the Betti number on different orders, a family of important topological properties used in biology [83, 84, 85] and physics [86, 87, 88] studies. In general, a $m$-order Betti number, $\mathcal{B}_{m}$, counts the number of $m$-dimensional holes in the system, which is a key topological characteristic of high-order structures and defines the dimension of $m$-th homology [74]. In the terminology of graph theory, a 00-order Betti number counts the number of connected components and a $1$-order Betti number is the number of cycles. Computationally, the $m$-order Betti number can be calculated using the toolbox offered by Ref.
@@ V.3 Identification of organizational structure
Figure 7: The organizational structure discovered by renormalization flows. (a-c) The Louvain communities underlying the Genetics (Ge), WI2007 (WI), and Inflamome (In) data sets [89] are extracted using the Louvain community detection algorithm [90]. The units of different communities are distinguished by colors. For comparison, the SRG is applied to renormalize these systems with $\left(p,q\right)=\left(1,2\right)$ and $\left(p,q\right)=\left(1,1\right)$. Units are determined as belonging to the same SRG community if they are aggregated into the same super-unit. Meanwhile, the corrected SRG communities are presented. (d-f) The correctness of communities and disconnectivity between communities are shown. (g-i) The adjusted mutual information, completeness, and homogeneity are measured on SRG communities. Notion “Ge (1,1)” denotes the SRG communities with $\left(p,q\right)=\left(1,1\right)$ on the Genetics data set. Notion “Ge Co” denotes the corrected SRG communities on the Genetics data set. Other notions can be understood in a similar way. Please see Appendix A for the definitions of all measures used in (d-i).
@@
When organizational structures are given, we explore the conditions under which these structures can be maintained by the SRG. In Appendix J, we implement the verification using protein-protein interactions (e.g., orthologous and genetic interactions) in Caenorhabditis elegans [89]. We first extract the initial community structures of protein-protein interactions by applying the Louvain community detection algorithm [90] (see Appendix A for introduction). Then, we use the SRG to renormalize protein-protein interactions on a high order (i.e., with $\left(p,q\right)=\left(1,2\right)$) or a low order (i.e., with $\left(p,q\right)=\left(1,1\right)$). We define latent community structures according to unit aggregation during renormalization (see Appendix J for details). Consistent with Ref. [38], our results in Appendix J highlight that renormalization can never be treated as a trivial counterpart of clustering. Meanwhile, the SRG guided by high-order interactions (e.g., $q=2$) can generally preserve the properties of initial community structures while the SRG guided by pairwise interactions does not. Based on these results, we speculate that high-order interactions are essential in forming and characterizing organizational structures. Communities can not be treated as the trivial collections of pairwise interactions and it may be inappropriate to partition a system into sub-systems only based on pairwise relations. Given this observation, we naturally wonder if high-order interactions alone are sufficient to define ideal community structures (i.e., do not consider pairwise relations while determining system partition).
@@
To tackle this question, we explore whether a SRG defined with different combination of $\left(p,q\right)$ can be applied to identify latent communities when system partition is unknown. In Fig. 7(a-c), we define communities according to the renormalization flows of the SRG and compare them with the Louvain communities (i.e., the communities detected by the Louvain algorithm [90], see Appendix A). Specifically, we classify the units into the same SRG community if they are aggregated into the same super-unit during renormalization (e.g., we generate SRG communities after the first iteration of renormalization flows in Fig. 7(a-c)).
@@ V.4 Optimization of information bottleneck
Finally, we analyze the SRG in terms of informational properties. In previous studies, the analysis using information theory has suggested that renormalization groups essentially maximize the mutual information between relevant features and the environment [94, 95] or reduce the mutual information among irrelevant features [96]. In this work, we suggest exploring the SRG from the perspective of information bottleneck [97] because renormalization groups, similar to dimensionality reduction or representation learning approaches in machine learning [98, 99, 100], essentially deal with information encoding during data compression [101].
Figure 8: The information curve of the SRG. (a) The SRG is applied on four synthetic interacting systems, whose pairwise interactions follow the Erdos-Renyi network (ER, each pair of units share an edge with a probability of $0.01$), the Watts-Strogatz network (WS, each unit initially has $5$ neighbors and edges are rewired according to a probability of $0.3$), the Barabási-Albert network (BA, $c=1$), and the random tree (RT), respectively. Each kind of system consists of $500$ units and has $10$ replicas. The SRG is defined with the multi-order Laplacian and $\left(p,q\right)=\left(1,1\right)$ for RT while $\left(p,q\right)\in\{\left(1,1\right),\left(1,2\right),\left(1,3\right)\}$ is set for all other systems. The generated information flows are used to derive information curves. (b) The complexity per preserved information bit is shown after averaging across all replicas.
@@
where $i=k-1$ if $k>1$ and $i=k$ if $k=1$. Parameter $\psi$ denotes the regularization strength. Notion $\mathcal{I}\left(\cdot;\cdot\right)$ denotes the mutual information (see Appendix A for its definition), which can be estimated by the mutual information estimator with local non-uniformity corrections (we set the correction strength as $10^{-3}$ such that corrections only happen when necessary) [103]. This approach has been included in the non-parametric entropy estimation toolbox [104]. Note that Eq. (31) is different from the information bottleneck considered in Ref. [101] due to the discrepancy between our concerned question and Ref. [101].
@@ VI Discussions
Various real interacting systems share universal characteristics that govern their dynamics and phase transition behaviours [105]. These characteristics remain elusive because fundamental physics tools, such as path integrals and renormalization group theories, are not well-established when a priori knowledge about system mechanisms is absent [10]. This gap may be accountable for diverse controversies concerning the critical phenomena in biological or social systems (e.g., the controversies about brain criticality [106]) and has induced a booming field that is devoted to developing statistical physics approaches for the empirical data generated by unknown mechanisms [10]. Notable progress is accomplished in the computational implementations of path integrals and renormalization groups on real systems with pairwise interactions [29, 30, 10, 31, 32, 33], supporting remarkable applications in analyzing neural dynamics [107] and swarm behaviours [108]. Nevertheless, there exists no apparent equivalence between these pioneering works and an appropriate approach for analyzing un-decomposable high-order interactions because classic network or geometric representations are invalid in characterizing polyadic relations [11, 12, 13, 9, 14].
In this work, we contribute to the field by developing the natural generalizations of path integrals and renormalization groups on un-decomposable high-order interactions. Our main contributions are summarized below:
@@
    We suggest that the contributions of microscopic fluctuations to macroscopic states in the moment space are precisely characterized by a function of the multi-order Laplacian [9] or our propsed high-order path Laplacian. Consequently, a renormalization group, the SRG, can be directly developed based on the diffusion on simplicial complexes, which does not require any assumption on latent metric for mapping units into target spaces (e.g., the hyperbolic space [30, 33]). This property ensures the general applicability of our theory on real complex systems, where a priori knowledge is rare and any assumption can be unreliable. Different from classic renormalization groups, the SRG can renormalize $p$-order interactions in a system based on the structure and dynamics associated with $q$-order interactions ($p\leq q$), enabling us to analyze the effects of $q$-order interactions on $p$-order ones and study the scaling properties across different orders.
@@ Acknowledgements
This project is supported by the Artificial and General Intelligence Research Program of Guo Qiang Research Institute at Tsinghua University (2020GQG1017) as well as the Tsinghua University Initiative Scientific Research Program. Authors appreciate Hedong Hou, who studies at the Institut de Mathématiques d’Orsay, for his inspiring discussions.
@@ Appendix A Auxiliary concept definitions
    The Erdos-Renyi network (ER) [81] is an elementary type of random network in which two units share an edge with a probability of $p$ or are disconnected with a probability of $1-p$. We have used the ER networks in Fig. 5, Fig. 8, and Figs. 10-11.
@@
    The Watts-Strogatz network (WS) [81] is a typical model of small-world phenomenon. It is generated by an edge rewiring process, during which we first initialize a regular network (i.e., a network where all units have the same number of neighbors) and rewire each edge according to a probability of $p$ (i.e., the edge remains unchanged with a probability of $1-p$). Several WS networks are implemented in Figs. 4-5, Fig. 8, and Figs. 9-11.
@@
    The Barabási-Albert network (BA) [47, 48] is a representative example of weakly scale-invariant system [10]. It is generated by the network growth process (i.e., a process for progressively adding new units and edges into a null network), where the number of random edges to attach from a new unit to existing units is denoted as $c$. One can see the BA networks in Figs. 1-3, Fig. 5, Fig. 8, and Figs. 10-11.
@@
    The random tree (RT) [111] is an instance of scale-invariance, which is generated from the uniformly random Prüfer sequence in computational implementation. The instances of RT are implemented in Figs. 4-5, Fig. 8, and Figs. 10-11.
@@
    The triangular lattice (TL) [112] is another instance of scale-invariance when system size approaches to infinity. Under the periodic boundary condition, a finite triangular lattice system is also a regular network where each unit has six neighbors. When the periodic boundary condition is absent, only the units within boundary form a regular network while the units on boundary may exhibit different behaviours under finite size effects. One can see the instances of TL in Figs. 4-5 and Figs. 10-11.
@@
    The square lattice (SL) [113] is also an instance of scale-invariance when system size approaches to infinity. It shares similar properties with the triangular lattice except that each unit in a finite square lattice system has four rather than six neighbors under the the periodic boundary condition. The SL has been analyzed in Fig. 4.
@@
In our experiments, we tackle the identification and preserving tasks of organizational structures of interacting systems, where we compare our SRG framework with the Louvain community detection algorithm [90, 92, 93]. The Louvain community detection algorithm searches for an ideal partition by maximizing the modularity of communities, which is essentially different from renormalization groups because it is purely optimization-based [90, 92, 93]. For convenience, we refer to the system partition identified by the Louvain community detection algorithm [90, 92, 93] as the Louvain communities.
When an organizational structure has been given (e.g., if the Louvain communities have been derived), we can explore whether the SRG can preserve this organizational structure during renormalization. The corresponding experiment is shown in Fig. 12 and Appendix J. We do not repeatedly elaborate the experiment in this section.
@@
    The adjusted mutual information [114] is a mutual information score (i.e., similarity degree) between the SRG and the Louvain communities, which is regulated by chance and is independent of permutation [114]. The adjusted mutual information approaches to $1$ when the compared communities are identical.
@@
    The completeness score [115] measures the consistency between the SRG and the Louvain communities, which approaches to $1$ if all units within each SRG community also belong to the same Louvain community.
@@
    The homogeneity score [115], similar to the completeness score, measures the consistency between the SRG and the Louvain communities, which approaches to $1$ if all units within each Louvain community also belong to the same SRG community.

### v7

@@ V.2 Exploration of the topological invariance
In Appendix C, we have explained how the SRG is theoretically related to the persistent homology theory and other concepts, where we suggest the potential applicability of the SRG in analyzing persistent homology. Below, we validate this applicability by using the SRG to explore the topological invariance.
In brief, the persistent homology theory can be used to study the lifetimes of different topological properties in a filtration process, during which we progressively add or remove simplices from the system [62, 74, 75, 76] (i.e., this is similar to the network evolution process in complex network studies [81]). Although the SRG is not exactly same as the classic filtration approach because it modifies the system in a more complicated manner, it does offer an opportunity to study the evolution of topological properties during renormalization (i.e., we can measure different topological properties on $\mathbf{G}^{\left(p\right)}_{k}$ for each $k\geq 1$). Similar to the analysis of scale-invariance, a topological property can be treated as invariant if it is persistent across all the iterations of the SRG (i.e., referred to as the topological invariance). In our analysis, we mainly focus on the Betti number on different orders, a family of important topological properties used in biology [82, 83, 84] and physics [85, 86, 87] studies. In general, a $m$-order Betti number, $\mathcal{B}_{m}$, counts the number of $m$-dimensional holes in the system, which is a key topological characteristic of high-order structures and defines the dimension of $m$-th homology [74]. In the terminology of graph theory, a 00-order Betti number counts the number of connected components and a $1$-order Betti number is the number of cycles. Computationally, the $m$-order Betti number can be calculated using the toolbox offered by Ref. [88].
@@ V.3 Identification of organizational structure
When organizational structures are given, we explore the conditions under which these structures can be maintained by the SRG. In Appendix L, we implement the verification using protein-protein interactions (e.g., orthologous and genetic interactions) in Caenorhabditis elegans [89]. We first extract the initial community structures of protein-protein interactions by applying the Louvain community detection algorithm [90] (see Appendix A for introduction). Then, we use the SRG to renormalize protein-protein interactions on a high order (i.e., with $\left(p,q\right)=\left(1,2\right)$) or a low order (i.e., with $\left(p,q\right)=\left(1,1\right)$). We define latent community structures according to unit aggregation during renormalization (see Appendix L for details). Consistent with Ref. [38], our results in Appendix L highlight that renormalization can never be treated as a trivial counterpart of clustering. Meanwhile, the SRG guided by high-order interactions (e.g., $q=2$) can generally preserve the properties of initial community structures while the SRG guided by pairwise interactions does not. Based on these results, we speculate that high-order interactions are essential in forming and characterizing organizational structures. Communities can not be treated as the trivial collections of pairwise interactions and it may be inappropriate to partition a system into sub-systems only based on pairwise relations. Given this observation, we naturally wonder if high-order interactions alone are sufficient to define ideal community structures (i.e., do not consider pairwise relations while determining system partition).
Figure 8: The organizational structure discovered by renormalization flows. (a-c) The Louvain communities underlying the Genetics (Ge), WI2007 (WI), and Inflamome (In) data sets [89] are extracted using the Louvain community detection algorithm [90]. The units of different communities are distinguished by colors. For comparison, the SRG is applied to renormalize these systems with $\left(p,q\right)=\left(1,2\right)$ and $\left(p,q\right)=\left(1,1\right)$. Units are determined as belonging to the same SRG community if they are aggregated into the same macro-unit. Meanwhile, the corrected SRG communities are presented. (d-f) The correctness of communities and disconnectivity between communities are shown. (g-i) The adjusted mutual information, completeness, and homogeneity are measured on SRG communities. “Ge (1,1)” denotes the SRG communities with $\left(p,q\right)=\left(1,1\right)$ on the Genetics data set. “Ge Co” denotes the corrected SRG communities on the Genetics data set. Other notions can be understood in a similar way. Please see Appendix A for the definitions of all measures used in (d-i).
@@
To tackle this question, we explore whether a SRG defined with different combination of $\left(p,q\right)$ can be applied to identify latent communities when system partition is unknown. In Fig. 8(a-c), we define communities according to the renormalization flows of the SRG and compare them with the Louvain communities (i.e., the communities detected by the Louvain algorithm [90], see Appendix A). Specifically, we classify the units into the same SRG community if they are aggregated into the same macro-unit during renormalization (e.g., we generate SRG communities after the first iteration of renormalization flows in Fig. 8(a-c)).
@@ V.4 Optimization of information bottleneck
Finally, we analyze the SRG in terms of informational properties. In previous studies, the analysis using information theory has suggested that renormalization group frameworks essentially maximize the mutual information between relevant features and the environment [94, 95] or reduce the mutual information among irrelevant features [96]. In this work, we suggest exploring the SRG from the perspective of information bottleneck [97] because a renormalization group, similar to dimensionality reduction or representation learning approaches in machine learning [98, 99, 100], essentially deals with information encoding during data compression [101].
@@
where $i=k-1$ if $k>1$ and $i=k$ if $k=1$. The parameter $\psi$ defines the regularization strength. We denote $\mathcal{I}\left(\cdot;\cdot\right)$ as the mutual information (see Appendix A for its definition), which can be estimated by the mutual information estimator with local non-uniformity corrections (we set the correction strength as $10^{-3}$ such that corrections only happen when necessary) [103]. This approach has been included in the non-parametric entropy estimation toolbox [104]. Note that Eq. (33) is different from the information bottleneck considered in Ref. [101] due to the discrepancy between our concerned question and Ref. [101].
Figure 9: The information curve of the SRG. (a) The SRG is applied on four synthetic interacting systems, whose pairwise interactions follow the Erdos-Renyi network (ER, each pair of units share an edge with a probability of $0.02$), the Watts-Strogatz network (WS, each unit initially has $5$ neighbors and edges are rewired according to a probability of $0.05$), the Barabási-Albert network (BA, $c=4$), and the random tree (RT), respectively. Each kind of system consists of $500$ units and has $5$ replicas. The SRG is defined with the multi-order Laplacian and $\left(p,q\right)=\left(1,1\right)$ for RT while $\left(p,q\right)\in\{\left(1,1\right),\left(1,2\right),\left(1,3\right)\}$ is set for all other systems. The generated information flows are used to derive information curves. (b) The complexity per preserved information bit is shown after averaging across all replicas.
@@ VI Discussions
Various real interacting systems share universal characteristics that govern their dynamics and phase transition behaviours [105]. These characteristics remain elusive because fundamental physics tools, such as path integrals and renormalization group theories, are not well-established when a priori knowledge about system mechanisms is absent [10]. This gap may be accountable for diverse controversies concerning the critical phenomena in biological or social systems (e.g., the controversies about brain criticality [106]) and has induced a booming field that is devoted to developing statistical physics approaches for the empirical data generated by unknown mechanisms [10]. Notable progress is accomplished in the computational implementations of path integral and renormalization group theories on real systems with pairwise interactions [29, 30, 10, 31, 32, 33], supporting remarkable applications in analyzing neural dynamics [107] and swarm behaviours [108]. Nevertheless, there exists no apparent equivalence between these pioneering works and an appropriate approach for analyzing un-decomposable high-order interactions because classic network or geometric representations are invalid in characterizing polyadic relations [11, 12, 13, 9, 14].
In this work, we contribute to the field by developing the natural generalizations of path integral and renormalization group on un-decomposable high-order interactions. Our main contributions are summarized below:
@@
    We suggest that the contributions of microscopic fluctuations to macroscopic states in the moment space are precisely characterized by a function of the multi-order Laplacian [9] or our propsed high-order path Laplacian. Consequently, a renormalization group, the SRG, can be directly developed based on the diffusion on simplicial complexes, which does not require any assumption on latent metric for mapping units into target spaces (e.g., the hyperbolic space [30, 33]). This property ensures the general applicability of our theory on real complex systems, where a priori knowledge is rare and any assumption can be unreliable. Different from classic renormalization groups, the SRG can renormalize $p$-order interactions in a system based on the structure and dynamics associated with $q$-order interactions ($p\leq q$), enabling us to analyze the effects of $q$-order interactions on the $p$-order ones and study the scaling properties across different orders.
@@ Acknowledgements
This work is a part of the Topophy program. Y.T develops the theory, designs computational experiments, writes the manuscript, and leads the Topophy program. A.H.C contributes to theoretical derivations, programmatic implementation, and computational experiment realization. Y.H.X contributes to theoretical derivations and proof reading. P.S supervises the research project, proofreads and revises the manuscript, contributes to conceptualization, and offers technical supports. This project is supported by the Artificial and General Intelligence Research Program of Guo Qiang Research Institute at Tsinghua University (2020GQG1017). Authors appreciate Dr. Hedong Hou, who studies at the Institut de Mathématiques d’Orsay, for his discussions. Authors thank Mr. Kangyu Weng, who studies at Tsinghua University, for his help in proof reading. Authors also appreciate anonymous reviewers for their inspiring suggestions, especially those about the topological invariance analysis, the SRG in a form of conventional renormalization group theories, and the nature of high-order interactions.
@@ Appendix A Auxiliary concept definitions
    The Erdos-Renyi network (ER) [80] is an elementary type of random network in which two units share an edge with a probability of $p$ or are disconnected with a probability of $1-p$. We have used the ER networks in Fig. 6, Fig. 9, Fig. 10, and Figs. 12-13.
@@
    The Watts-Strogatz network (WS) [111] is a typical model of small-world phenomenon. It is generated by an edge rewiring process, during which we first initialize a regular network (i.e., a network where all units have the same number of neighbors) and rewire each edge according to a probability of $p$ (i.e., the edge remains unchanged with a probability of $1-p$). Several WS networks are implemented in Figs. 5-6, Fig. 9, Fig. 10, and Figs. 11-13.
@@
    The Barabási-Albert network (BA) [47, 48] is a representative example of weakly scale-invariant system [10]. It is generated by the network growth process (i.e., a process for progressively adding new units and edges into a null network), where the number of random edges to attach from a new unit to existing units is denoted as $c$. One can see the BA networks in Figs. 1-3, Fig. 6, Fig. 9, Fig. 10, and Figs. 12-13.
@@
    The random tree (RT) [112] is an instance of scale-invariance, which is generated from the uniformly random Prüfer sequence in computational implementation. The instances of RT are implemented in Figs. 5-6, Fig. 9, Fig. 10, and Figs. 12-13.
@@
    The triangular lattice (TL) [113] is another instance of scale-invariance when system size approaches infinity. Under the periodic boundary condition, a finite triangular lattice system is also a regular network where each unit has six neighbors. When the periodic boundary condition is absent, only the units within boundary form a regular network while the units on boundary may exhibit different behaviours under finite size effects. One can see the instances of TL in Figs. 5-6, Fig. 10, and Figs. 12-13.
@@
    The square lattice (SL) [114] is also an instance of scale-invariance when system size approaches infinity. It shares similar properties with the triangular lattice except that each unit in a finite square lattice system has four rather than six neighbors under the the periodic boundary condition. The SL has been analyzed in Fig. 5 and Fig. 10.
@@
In our experiments, we tackle the identification and preserving tasks of organizational structures of interacting systems, where we compare our SRG framework with the Louvain community detection algorithm [90, 92, 93]. The Louvain community detection algorithm searches for an ideal partition by maximizing the modularity of communities, which is essentially different from a renormalization group because it is purely optimization-based [90, 92, 93]. For convenience, we refer to the system partition identified by the Louvain community detection algorithm [90, 92, 93] as the Louvain communities.
When an organizational structure has been given (e.g., if the Louvain communities have been derived), we can explore whether the SRG can preserve this organizational structure during renormalization. The corresponding experiment is shown in Fig. 14 and Appendix L. We do not repeatedly elaborate the experiment in this section.
@@
    The adjusted mutual information [115] is a mutual information score (i.e., similarity degree) between the SRG and the Louvain communities, which is regulated by chance and is independent of permutation [115]. The adjusted mutual information approaches $1$ when the compared communities are identical.
@@
    The completeness score [116] measures the consistency between the SRG and the Louvain communities, which approaches $1$ if all units within each SRG community also belong to the same Louvain community.
@@
    The homogeneity score [116], similar to the completeness score, measures the consistency between the SRG and the Louvain communities, which approaches $1$ if all units within each Louvain community also belong to the same SRG community.